OpenStudy (anonymous):

If f(x) = 3x + 1 and g(x) = 2x − 4, find f[g(2)]. 25 4 0 1

OpenStudy (anonymous):

u were given explanation by @whpalmer4 ur last sum its the same methode try it .... give ur ans ....

OpenStudy (jhannybean):

Ah..

OpenStudy (jhannybean):

First find f(g(x))... you have your two functions, plug in g(x) in place of "x" in your function f(x) \[\large f(g(x))= 3(2x-4)+1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 25

OpenStudy (anonymous):

@Jhannybean .... first substitute in g(x) the value of 2 then u will get the value of g(2).... substitute that in f(x)....

OpenStudy (anonymous):

@Trai7

OpenStudy (jhannybean):

Ohh, i forgot the 2 sub. haha nvm! i got it :D

OpenStudy (anonymous):

g(2)=2*2-4=0 f(x)=(3x+1) f(g(2))=3*0+1=0+1=1 ur ans is 1.... clear ...@Trai7

OpenStudy (anonymous):

@Trai7